40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Feb 7 15:00:16 2020, gautam, Update, LSC, More high BW POY experiments AO_TFs.pdf
    Reply  Fri Feb 7 19:39:10 2020, Koji, Update, LSC, More high BW POY experiments 
Message ID: 15200     Entry time: Fri Feb 7 19:39:10 2020     In reply to: 15199
Author: Koji 
Type: Update 
Category: LSC 
Subject: More high BW POY experiments 

This measurement tells you how the gain balance between the SLOW_CM and AO paths should be. Basically, what you need is to adjust the overall gain before the branch of the paths.

Except for the presence of the additional pole-zero in the optical gain because of the power recycling.

You have compensated this with a filter (z=120Hz, p=5kHz) for the CM path. However, AO path still don't know about it. Does this change the behavior of the cross over?

If the servo is not unconditionally stable when the AO gain is set low, can we just turn on the AO path at the nominal gain? This causes some glitch but if the servo is stable, you have a chance to recover the CARM control before everything explodes, maybe?

ELOG V3.1.3-